OpenStudy (anonymous):

I have $1.10 worth of coins in my pocket. The only types of coins are nickles and dimes. If there are 16 coins how many of each coin are there

OpenStudy (johnweldon1993):

this is a system of equations only nickels and dimes n + d = 16 .05n + .10d = 1.10 can you solve this?

OpenStudy (whpalmer4):

I like doing coin problems in cents, but otherwise do it exactly as @johnweldon1993 has done here: n + d = 16 5n + 10d = 110 The advantage to using cents instead of dollars is that you don't have to deal with decimals — but you should get the same answer either way.

OpenStudy (johnweldon1993):

yeah @whpalmer4 I was thinking about doing it without decimals...more fun this way though! more chances to make mistakes so better to concentrate :D lol

OpenStudy (johnweldon1993):

you with us @fridayman ?

OpenStudy (whpalmer4):

it's funny, in other pursuits, I'm a big fan of doing things the less comfortable way, so that when forced to do it the less comfortable way, it isn't so unfamiliar or difficult. In math, however, I seem not to do that :-)

OpenStudy (anonymous):

Yes, so once you get n + d = 10 .05n + .10d = 1.10 Is it just basically a substitution problem from there?

OpenStudy (whpalmer4):

I know some kids who do these problems like this: you've got 16 coins. Let them all be nickels. Then you would have 16 * 5 cents = 80 cents. That's not enough, so let's exchange some of those nickels for dimes. Each one you change bumps the total value up by 10-5 = 5 cents. 80 cents was $1.10 - 80 cents = 30 cents too low. 30 cents / 5 cents per exchange = 6 coins to exchange. That means 10 nickels and 6 dimes...

OpenStudy (johnweldon1993):

Basically yes...so since you seem to know what to do....just let us know what you get for an answer

OpenStudy (johnweldon1993):

@whpalmer4 O.o oh no! haha that's too much XD perhaps I wouldn't go THAT complicated! haha

OpenStudy (whpalmer4):

My grandmother kept the books for the family business, and used to accumulate huge amounts of change, so we did problems like this all the time when I was a little kid :-)

OpenStudy (johnweldon1993):

@whpalmer4 See...that's an awesome way to get started into math! haha kudos to your grandmother!

OpenStudy (whpalmer4):

any kid gets interested in arithmetic when there's money involved :-)

OpenStudy (johnweldon1993):

this is true! haha...actually it's funny how many kids I tutor...when I change it into a money problem or a "friends" problem...they'll get it like *snap* that lol...but then put it in mathematical terms and the looks would be O.o haha

OpenStudy (anonymous):

I cant get it

OpenStudy (johnweldon1993):

well lets see what you are doing...post what steps you were doing...

OpenStudy (anonymous):

0.5(n + d = 10) 0.5n + .10d = 1.10 0.5n + 0.5d = 5 0.5n + .10d = 1.10

OpenStudy (anonymous):

I get -0.5d = 3.9

OpenStudy (texaschic101):

n + d = 16 ---> n = 16 - d now sub 10 - d in for n in the other equation .05n + .10d = 1.10 .05(16 - d) + .10d = 1.10 .80 - .05d + 10d = 1.10 .05d = 1.10 - .80 .05d = .30 d = .30/.05 d = 6 now sub 6 in for d n + d = 16 n + 6 = 16 n = 16 - 6 n = 10 check.... .05n + .10d = 1.10 .05(10) + .10(6) = 1.10 .50 + .60 = 1.10 1.10 = 1.10 (correct) ANSWER : There are 10 nickels totaling 50 cents and 6 dimes totaling 60 cents. All together, there are 16 coins totaling 1.10. Any questions ?

OpenStudy (whpalmer4):

@fridayman where were you getting 0.5 from? a nickel is $0.05, or 5 cents. @texaschic101 showed you substitution, I'll show you elimination: \[n+d = 16\]\[5n+10d = 110\] Multiply first equation by -5: \[-5n-5d=-80\]\[5n+10d=110\]------------------ adding \[0n + 5d = 3\]\[5d=30\]\[ d=6\]\[n+d=16\]\[n+6=16\]\[n=10\] checking: \[5(10) + 10(6) = 50+60 = 110\checkmark\]